# Towards Safer Pretraining: Analyzing and Filtering Harmful Content in Webscale datasets for Responsible LLMs

### v1

@@ Abstract
Large language models (LLMs) have become integral to various real-world applications, leveraging massive, web-sourced datasets like Common Crawl, C4, and FineWeb for pretraining. While these datasets provide linguistic data essential for high-quality natural language generation, they often contain harmful content, such as hate speech, misinformation, and biased narratives. Training LLMs on such unfiltered data risks perpetuating toxic behaviors, spreading misinformation, and amplifying societal biases which can undermine trust in LLM-driven applications and raise ethical concerns about their use. This paper presents a large-scale analysis of inappropriate content across these datasets, offering a comprehensive taxonomy that categorizes harmful webpages into Topical and Toxic based on their intent. We also introduce a prompt evaluation dataset, a high-accuracy Topical and Toxic Prompt (TTP), and a transformer-based model (HarmFormer) for content filtering. Additionally, we create a new multi-harm open-ended toxicity benchmark (HAVOC) and provide crucial insights into how models respond to adversarial toxic inputs. Upon publishing, we will also opensource our model signal on the entire C4 dataset. Our work offers insights into ensuring safer LLM pretraining and serves as a resource for Responsible AI (RAI) compliance.
@@ 1 Introduction
Large language models (LLMs) Brown (2020), Achiam et al. (2023), Bubeck et al. (2023), Abdin et al. (2024), Dubey et al. (2024) derive their capabilities from pretraining on massive web-scale datasets such as Common CrawlRana (2010), C4 Raffel et al. (2020), and FineWeb Penedo et al. (2024a). These datasets, while diverse, contain significant amounts of harmful content, hate speech, misinformation, and explicit material, which propagate into LLM output. For instance, as shown in Figure 1, prompting a Qwen2 model to generate self-harm methods yields dangerous, actionable suggestions. To mitigate such risks, practitioners rely on content moderation tools like keyword blocklists [^1] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words, HateBERT Caselli et al. (2021), and LlamaGuard Inan et al. (2023). However, these tools suffer from critical limitations that undermine their effectiveness in real-world scenarios.
@@
To address these limitations, we propose a three-dimensional taxonomy (Safe, Topical, Toxic) that classifies text by their intent and severity across five harm categories, including Hate & Violence, Ideological Harm, Sexual Harm, Self-Inflicted Harm, and Illegal Activity Harm. Unlike binary frameworks, our taxonomy distinguishes harmful intent (e.g., promoting self-harm) from critical topical discourse (e.g., mental health resources), enabling nuanced content moderation. We operationalize this taxonomy through two scalable solutions: (1) TTP, a prompt-based on OpenAI’s GPT4-Omni, and (2) a LongFormer-based model Beltagy et al. (2020) optimized for long-form content moderation. Our pipeline processes web-scale datasets like Common Crawl, C4, and FineWeb, where manual curation is infeasible due to their trillion-token size. To support reproducibility, we will release TTP-Eval, a benchmark of annotated web pages spanning diverse harms and document lengths, enabling systematic evaluation of long-text moderation tools.
Existing benchmarks like RealToxicityPrompts (RTP) Gehman et al. (2020) focus narrowly on toxicity, overlooking nuanced harms such as conspiracy theories or biased narratives. We introduce HAVOC, a multidimensional benchmark that evaluates LLM safety across our taxonomy’s dimensions. HAVOC measures how language models generate harmful responses even to seemingly benign inputs, such as prompts about “vaccine efficacy” that elicit misinformation. By analyzing 10376 open-ended generations, we demonstrate that 26.7% of outputs from state-of-the-art LLMs exhibit harmful intent under our taxonomy. These findings underscore the limitations of current safety frameworks and motivate our contributions, which are as follows:
@@
  • Moderation Tools: We release TTP-Eval, an annotated benchmark for long-text moderation; TTP, a prompt-based classifier; and a LongFormer-based model that filters harmful content with high accuracy while preserving topical text. We use the model to filter toxic content from C4 and FineWeb datasets.
  • HAVOC Benchmark: A multidimensional multiharm benchmark to evaluate LLM safety, systematically exposing vulnerabilities in pre-trained language models over certain harms.
@@ 2 Related Works
Dataset Analysis and Filtering: Large pretraining corpora like Common Crawl Rana (2010), C4 Raffel et al. (2020), and FineWeb Penedo et al. (2024a) underpin modern NLP systems, but their unfiltered nature raises safety concerns. Luccioni and Viviano (2021) analyzed sexual and hateful content in Common Crawl using keyword heuristics, while Jansen et al. (2022) proposed perplexity-based filtering to remove toxic text. These methods lack precision: keyword filters conflate harmful intent with educational content (e.g., medical discussions of self-harm), and perplexity thresholds fail to capture contextual nuances. Recent studies caution against indiscriminate filtering, as overly aggressive removal may inadvertently amplify toxicity in model outputs Longpre et al. (2024).
Toxicity Evaluation Benchmarks: Efforts to evaluate LLM safety in open ended generation include RealToxicityPrompts (RTP) Gehman et al. (2020), a benchmark of 100K prompts using Perspective API to measure toxicity in open-ended generation. However, RTP focuses narrowly on hate and sexual content, omitting harms like Ideological Harms and Self-Inflicted Harms. Perspective API’s classification also struggles with adversarial cases, such as topical prompts that elicit harmful outputs, and lacks intent-aware evaluation.
Existing tools and benchmarks face three limitations: (1) oversimplification of harm detection to binary classification, (2) sentence-level focus ill-suited for long-form web text, and (3) narrow coverage of harm categories. Our work addresses these gaps through a three-dimensional taxonomy (Safe, Topical, Toxic) and a suite of tools for intent-aware moderation at scale.
@@ 3 Three-Dimensional Safety Taxonomy for LLM Risk Mitigation
Modern LLM pretraining relies on web-scale datasets whose unfiltered nature introduces critical safety risks. As detailed in Table 1, we propose a three-dimensional classification framework that addresses these risks through granular content analysis across five fundamental harm categories. This taxonomy enables systematic identification of harmful content while preserving contextually valuable material essential for model capability.
@@ 4.2 TTP Prompt Creation
We developed our prompt (TTP) through an iterative process to improve the performance on the prompt eval dataset (TTP-Eval). The prompt operates on a set of clear guidelines to ensure accurate classification, focusing on whether the thin line between simply discussing or actively promoting harmful content. The results yield a structured output in the form of a JSON object for easy integration into further analysis. Few-Shot prompting Brown et al. (2020), Chain of Thought Prompting Wei et al. (2023) helped us get better quality on the held out development set. The prompt’s performance on the test set is summarized in Table 3. The lower precision and recall of the prompt on Hate & Violence is explained by the ”Reporting webpages”. A reporting web page is a webpage purely reporting an incident of hate, or violence or any harm, with no intent of justifying or promoting the toxiciy included with the incident. These are those web pages that purely report hateful statements from a source against a receiver, with the source and the receiver being a person, organization or an entity.
@@
We benchmark Perspective API on our TTP-Evalset in table 4 to compare and highlight the quality and coverage of TTP prompt. Since the Perspective API is predominantly trained for hate, sexual and bias, it exhibits lower quality on our TTP-Evalset. To cope for the text length difference between our set and the way the API is trained, we chunk the texts into a length of 500 characters, and take the maximum score across chunks with 0.4 as the optimal threshold chosen from the dev set.
@@ 5.1 Data Creation
We utilized the high-quality TTP, to annotate a dataset comprising 3 million web pages from various sources, including Common Crawl (1 Million), C4 Dodge et al. (2021) (1 Million) and FineWeb Penedo et al. (2024b) (1 Million). From this corpus, we select 258,000 documents, sampling based on both the source—Common Crawl, C4, and FineWeb—and the associated harms and topics. To ensure a balanced representation across all classes, we undersampled the majority of safe categorized web pages from the original 3 million prompt candidates. The labeled data was then stratified sampled and partitioned into three subsets for training and testing the model: training, development, and test sets, with a 90:5:5 split. This gives us a training set of 253,000 samples, and development and test sets of 14,000 each.
@@ 5.2 Model Training
We built HarmFormer, leveraging several Transformer Vaswani et al. (2023) based models - XLM-Roberta Conneau et al. (2020), Hate-BERT Caselli et al. (2021) and Longformer Beltagy et al. (2020). Our model utilizes the body of the text as its input, to assign labels across harms and dimensions. We employ a Multi-Task architecture with five distinct classification heads representing the harm categories, where each harm category head again categorizes the input into one of three dimensions: Safe, Topical, or Toxic. This design captures the nuanced boundaries between the Topical and Toxic dimensions within each harm category, as well as between the Safe and Topical dimensions, allowing for more precise classification across the task.
@@
The detailed results of our experiments are presented in Table 5. As expected, Hate-BERT Caselli et al. (2021), which is primarily trained to detect hate speech, did not perform well on our test set (at a tuned threshold of 0.05). When finetuned to our scenario, Hate-BERT still underperforms with other models as it needs to relearn new harm categories along with the dimensions. We leverage Long Former’s (149M Parameters) 1024 context length capability over XLM-Roberta, which gave us an F1-Gain of 2 points over the same model trained with 512 context length. We finalize the architecture of the our HarmFormer model to be a LongFormer Beltagy et al. (2020), trained with 1024 context length across layers, a batch size of 16, learning rate of 2e-5, with AdamW Loshchilov and Hutter (2019) as the optimizer for 3 epochs over the train set.
@@ 6.3 HAVOC: Harmful Abstractions and Violations in Open Completions Benchmark
Contrary to Americans, exactly who are more acquainted with showing all their emotions through consumer displays, Western european ladies opt to wait until they are simply alone to talk about what they think about you or your actions. Topical This is because European women of all ages have a tendency to be even more reserved and do not like to show their feelings in public areas. The best way to get acquainted with an American woman should be to ask her out on a date. Toxic
@@
We leverage Gemma 2 Gemma (2024), Llama 3.2 Meta (2024), and Mistral V0.3 Mistral (2024) family of models to study their open ended generation on HAVOC and RTP. We use the base pretrained models for our evaluations. Due to hardware constraints, we leverage 2B, 9B & 27B model variants from Gemma 2, 1B & 3B variants from Llama 3.2, and use the 7B variant from the Mistral V0.3 family. We use Ollama Ollama (2024) to generate the completions for all the Language Models on our benchmark. We use greedy decoding to generate 200 output tokens from each prefix.
@@ 7 Conclusion
Our findings reveal widespread harmful content in web-crawled datasets like Common Crawl, C4 and FineWeb. We also provide the major toxic topics in the harms we define, along with their proportions. Through a combination of manual annotation, prompt-based labeling, and transformer models, we demonstrate that our semi-automated approaches can scale to handle the challenge of inappropriate content detection across all harms. We provide a publicly available eval-set for web pages, a prompt to identify the same, transformer-based baseline model and its Topical and Toxic scores on C4 and FineWeb. We give analyze and quantify issues in open text datasets like Common Crawl, C4 and FineWeb by covering across various topics, harms, and dimensions. Our multi harm open ended toxic generation benchmark, HAVOC quantifies the toxicity in language model’s outputs when challenged with sensitive and provocative inputs. We also find sexual as the most leaking harm for a language model when provocated. Having highlighted the problem in the large scale text datasets, and pre-trained language models, we offer a valuable resource for developing content filtering methods and a new superior open ended text generation benchmark that quantifies toxicity in language models.
@@ 8 Limitations & Future Work
This study focuses on English data, leaving multilingual analysis for future work. The proposed taxonomy addresses broad harms but does not account for culture-specific sensitivities, such as actions deemed disrespectful in certain cultural contexts. For instance, a webpage encouraging head massages may be offensive in cultures where the head is considered sacred. Additionally, we aim to train a model on fully clean datasets to evaluate the impact of dataset purity on model behavior and performance, comparing these findings with models trained on unclean datasets. This will allow for a deeper understanding of the interplay between ”cleanliness” of training data and the model’s behaviour with toxic inputs.

### v2

@@ Abstract
Large language models (LLMs) have become integral to various real-world applications, leveraging massive, web-sourced datasets like Common Crawl, C4, and FineWeb for pretraining. While these datasets provide linguistic data essential for high-quality natural language generation, they often contain harmful content, such as hate speech, misinformation, and biased narratives. Training LLMs on such unfiltered data risks perpetuating toxic behaviors, spreading misinformation, and amplifying societal biases which can undermine trust in LLM-driven applications and raise ethical concerns about their use. This paper presents a large-scale analysis of inappropriate content across these datasets, offering a comprehensive taxonomy that categorizes harmful webpages into Topical and Toxic based on their intent. We also introduce a prompt evaluation dataset, a high-accuracy Topical and Toxic Prompt (TTP), and a transformer-based model (HarmFormer) for harmful content filtering. Additionally, we create a new multi-harm open-ended toxicity benchmark (HAVOC) and provide crucial insights into how models respond to adversarial toxic inputs. We share TTP, TTP-Eval, HAVOC and a sample of C4 inferenced on HarmFormer here  [^1] https://github.com/themendu/TowardsSaferPretraining. Our work offers insights into ensuring safer LLM pretraining and serves as a resource for Responsible AI (RAI) compliance.
@@ 1 Introduction
Large language models (LLMs) Achiam et al. (2023), Bubeck et al. (2023), Abdin et al. (2024), Dubey et al. (2024) derive their capabilities from pretraining on massive web-scale datasets such as Common CrawlRana (2010), C4 Raffel et al. (2020), and FineWeb Penedo et al. (2024). These datasets, while diverse, contain significant amounts of harmful content, hate speech, misinformation, and explicit material, which propagate into LLM output. For instance, as shown in Figure 1, prompting a Qwen2 model to generate self-harm methods yields dangerous, actionable suggestions. To mitigate such risks, practitioners rely on content moderation tools like keyword blocklists [^2] https://github.com/LDNOOBW/List-of-Dirty-Naughty-Obscene-and-Otherwise-Bad-Words, HateBERT Caselli et al. (2021), and Perspective API  [^3] https://perspectiveapi.com/. However, these tools suffer from critical limitations that undermine their effectiveness in real-world scenarios.
@@
To address these limitations, we propose a three-dimensional taxonomy (Safe, Topical, Toxic) that classifies text by their intent and severity across five harm categories, including Hate & Violence, Ideological Harm, Sexual Harm, Self-Inflicted Harm, and Illegal Activity Harm. Unlike binary frameworks, our taxonomy distinguishes harmful intent (e.g., promoting self-harm) from critical topical discourse (e.g., mental health resources), enabling nuanced content moderation. We operationalize this taxonomy through two scalable solutions: (1) TTP, a prompt-based on OpenAI’s GPT4-Omni, and (2) HarmFormer, a LongFormer-based model Beltagy et al. (2020) optimized for long-form content moderation. Our pipeline processes web-scale datasets like Common Crawl, C4, and FineWeb, where manual curation is infeasible due to their trillion-token size. To support reproducibility, we will release TTP-Eval, a benchmark of annotated web pages spanning diverse harms and document lengths, enabling systematic evaluation of long-text moderation tools.
Existing benchmarks like RealToxicityPrompts (RTP) Gehman et al. (2020) focus narrowly on toxicity, overlooking nuanced harms such as conspiracy theories or biased narratives. We introduce HAVOC, a multidimensional benchmark that evaluates LLM safety across our taxonomy’s dimensions. HAVOC measures how language models generate harmful responses even to seemingly benign inputs, such as prompts about “vaccine efficacy” that elicit misinformation. By analyzing 10376 open-ended generations, we demonstrate that 26.7% of outputs from state-of-the-art LLMs exhibit harmful content. These findings underscore the limitations of current safety frameworks and motivate our contributions, which are as follows:
@@
  • Moderation Tools: We release TTP-Eval, an annotated benchmark for long-text moderation; TTP, a prompt-based classifier; and HarmFormer, a LongFormer-based model that filters harmful content with high accuracy while preserving topical text.
  • HAVOC Benchmark: A multidimensional multiharm benchmark to evaluate LLM safety.
@@ 2 Related Works
Dataset Analysis and Filtering: Large pretraining corpora like Common Crawl Rana (2010), C4 Raffel et al. (2020), and FineWeb Penedo et al. (2024) underpin modern NLP systems, but their unfiltered nature raises safety concerns. Luccioni and Viviano (2021) analyzed sexual and hateful content in Common Crawl using keyword heuristics, while Jansen et al. (2022) proposed perplexity-based filtering to remove toxic text. These methods lack precision: keyword filters conflate harmful intent with educational content (e.g., medical discussions of self-harm), and perplexity thresholds fail to capture contextual nuances. Recent studies caution against indiscriminate filtering, as overly aggressive removal may inadvertently amplify toxicity in model outputs Longpre et al. (2024).
Toxicity Evaluation Benchmarks: Real Toxicity Prompts Gehman et al. (2020) (100K prompts + Perspective API scores) tests LLM safety, but its harm coverage (Hate and Sexual Content) is narrow and Perspective API can be misled by adversarial wording and misses intent and other harms.
Existing works face three limitations: (1) oversimplification of harm detection to binary classification, (2) short-text focus ill-suited for long-form web text, and (3) narrow coverage of harm categories. Our work addresses these gaps through a three-dimensional taxonomy (Safe, Topical, Toxic) and a suite of tools for intent-aware moderation at scale.
@@ 3 Three-Dimensional Safety Taxonomy for LLM Risk Mitigation
As detailed in Table 1, we propose a three-dimensional classification framework that addresses these risks through granular content analysis across five fundamental harm categories. This taxonomy enables systematic identification of harmful content while preserving contextually valuable material essential for model capability.
@@ 4.2 TTP Prompt Creation
We developed our prompt (TTP) through an iterative process to improve the performance on the prompt eval dataset (TTP-Eval) on GPT-4 Omni. The prompt operates on a set of clear guidelines to ensure accurate classification, focusing on whether the thin line between simply discussing or actively promoting harmful content. Few-Shot prompting Brown et al. (2020), Chain of Thought Prompting Wei et al. (2022) helped us get better quality on the held out development set. The prompt’s performance on the test set is summarized in Table 4. The lower precision and recall of the prompt on Hate & Violence is explained by the ”Reporting webpages”. A reporting web page is a webpage purely reporting an incident of hate, or violence or any harm, with no intent of justifying or promoting the toxiciy included with the incident. These are those web pages that purely report hateful statements from a source against a receiver, with the source and the receiver being a person, organization or an entity.
@@
We benchmark Perspective API on our TTP-Evalset in table 4 to compare and highlight the quality and coverage of TTP prompt. Since the Perspective API is predominantly trained for hate, sexual and bias, it exhibits lower quality on our TTP-Evalset. To cope for the higher text lengths in TTP-Evalset, we chunk the texts into a length of 500 characters, and take the maximum score across chunks with 0.4 as the optimal threshold chosen from the dev set. We also run TTP on R1 DeepSeek-AI et al. (2025), Gemma 2 Team et al. (2024) and Gemini 2, and as expected the models perform poorly, as the prompt is tuned with GPT 4 Omni.
@@ 5.1 Data Creation
We utilized the high-quality TTP, to annotate a dataset comprising 3 million web pages from various sources, including Common Crawl (1 Million), C4 Dodge et al. (2021) (1 Million) and FineWeb Penedo et al. (2024) (1 Million). From this corpus, we select 258,000 documents, sampling based on both the source—Common Crawl, C4, and FineWeb—and the associated harms and topics. To ensure a balanced representation across all classes, we undersampled the majority of safe categorized web pages from the original 3 million prompt candidates. The labeled data was then stratified sampled and partitioned into three subsets for training and testing the model: training, development, and test sets, with a 90:5:5 split. This gives us a training set of 253,000 samples, and development and test sets of 14,000 each.
@@ 5.2 Model Training
We built HarmFormer, leveraging several Transformer Vaswani et al. (2017) based models - XLM-Roberta Conneau et al. (2020), Hate-BERT and Longformer Beltagy et al. (2020). We employ a Multi-Task architecture with five distinct classification heads representing the harm categories, where each harm category head again categorizes the input into one of three dimensions: Safe, Topical, or Toxic with the web page’s text as the input. This design captures the nuanced boundaries between the Topical and Toxic dimensions within each harm category, as well as between the Safe and Topical dimensions, allowing for more precise classification across the task.
@@
The detailed results of our experiments are presented in Table 5. As expected, Hate-BERT, which is primarily trained to detect hate speech, did not perform well on our test set (at a tuned threshold of 0.05). When finetuned to our scenario, Hate-BERT still underperforms with other models as it needs to relearn new harm categories along with the dimensions. We leverage Long Former’s (149M Parameters) 1024 context length capability over XLM-Roberta, which gave us an F1-Gain of 2 points over the same model trained with 512 context length. We finalize the architecture of the our HarmFormer model to be a LongFormer Beltagy et al. (2020), trained with 1024 context length across layers, a batch size of 16, learning rate of 2e-5, with AdamW Loshchilov and Hutter (2019) as the optimizer for 3 epochs over the train set.
@@ 6.3 HAVOC: Harmful Abstractions and Violations in Open Completions Benchmark
In 1987, Dr. Hans Hertel of Germany published his findings that microwaves cause changes to DNA (the genetic material) which can lead to cancer. He also found that microwaved foods have a lower nutritional value than non-microwaved foods. Topical This is because European women of all ages have a tendency to be even more reserved and do not like to show their feelings in public areas. Toxic
@@
We leverage Gemma 2 Team et al. (2024), Llama 3.2 Meta (2024), and Mistral V0.3 Mistral (2024) family of models to study their open ended generation on HAVOC and RTP. We use the base pretrained models for our evaluations. Due to hardware constraints, we leverage 2B, 9B & 27B model variants from Gemma 2, 1B & 3B variants from Llama 3.2, and use the 7B variant from the Mistral V0.3 family. We use Ollama Ollama (2024) to generate the completions for all the Language Models on our benchmark. We use greedy decoding to generate 200 output tokens from each prefix.
@@ 7 Conclusion
Our findings reveal widespread harmful content in web-crawled datasets like Common Crawl, C4 and FineWeb. We also provide the major toxic topics in the harms we define, along with their proportions. Through a combination of manual annotation, prompt-based labeling, and transformer models, we demonstrate that our semi-automated approaches can scale to handle the challenge of inappropriate content detection across all harms. We provide a publicly available eval-set, prompt and a transformer based model for toxic text identification in web pages. We analyze and quantify issues in open text datasets like Common Crawl, C4 and FineWeb by covering across various topics, harms, and dimensions. Our multi harm open ended toxic generation benchmark, HAVOC quantifies the toxicity in language model’s outputs when challenged with sensitive and provocative inputs. We also find sexual as the most leaking harm for a language model when provocated. Having highlighted the problem in the large scale text datasets, and pre-trained language models, we offer a valuable resource for developing content filtering methods and a new superior open ended text generation benchmark that quantifies toxicity in language models.
@@ 8 Limitations & Future Work
This study focuses on English data, leaving multilingual analysis for future work. The proposed taxonomy addresses broad harms but does not account for culture-specific sensitivities, such as actions deemed disrespectful in certain cultural contexts. For instance, a webpage encouraging head massages may be offensive in cultures where the head is considered sacred. Additionally, we aim to train a model on dimensional filtered datasets and evaluate the impact of dataset purity on model behavior and performance, comparing these findings with models trained on unclean datasets. This will allow for a deeper understanding of the interplay between the cleanliness of training data and the model’s behaviour when faced with seemingly toxic inputs across multiple harms.